Fill the table

Refractive index (n)

Radius of curvature (R), m

Focal length (f), m

          1.2

          1.5

          1.5

          0.8

          1.0

          1.0

          1.6

          0.92

          1.8

          1.2
Answer #1

This is about Lendsmakers problem

This\:equation\:is\:defined\:as:

\frac{1}{f}=\left ( n-1 \right )\left ( \frac{1}{R_1}+\frac{1}{R_2} \right )

wherre

f-focal length

n=Refractive\:index

R_1=radius\:of\:curvature\:1

R_2=radius\:of\:curvature\:2

In\:this\:case\:\left ( by\:table \right )\:they\:consider\:just\:1\:radius\:of\:curvature.

Substituting\:the\:values\:from\:table:

1:

R=\left ( n-1 \right )f

R=\left ( 1.2-1 \right )1.5={\color{Blue} 0.3\:m}

2:

f=\frac{R}{\left ( n-1 \right )}={\color{Blue} 1.6\:m}

3:

n=\frac{R}{f}+1={\color{Blue} 2}

4:

R=\left ( n-1 \right )f={\color{Blue} 0.552\:m}

5:

f=\frac{R}{\left ( n-1 \right )}={\color{Blue} 1.5\:m}